# Topological constraints on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations

### v2

@@ 3.2.3 Application of $\mathbb{Q}$-valued augmentations
Figure 2 depicts $u(\mathbb{R}\times[0,1])\subset T^{*}\mathbb{R}^{3}$ and the three paths in $L_{K_{0}}$. The group $\mathbb{R}$ acts on $\mathcal{M}_{(H,J),A}(x,y)$ by $(u\cdot\sigma)(s,t)=u(s+\sigma,t)$ for every $\sigma\in\mathbb{R}$.
To define the $\mathbb{Z}/2$-graded Floer cochain complex, we refer to [18, Section 2.f]. We take the pair $(H,J)$ to be regular so that the moduli space $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is cut out transversely for every $x,y\in P_{H}\cap L_{K_{0}}$ such that $x\neq y$. The dimension of $\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is equal to |x||y|1|x|-|y|-1 modulo $2$. Moreover, by the Gromov compactness, the union of 0-dimensional components of $\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is a finite set. We define $n_{(H,J),A}(x,y)\in\mathbb{F}_{2}$ to be the modulo $2$ count of the number of the 0-dimensional components of $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$.
@@
For $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ in the statement of Proposition 2.7, let us set $\psi\coloneqq\varphi^{-1}$. Then, we can change the setup as follows: $\psi(\mathbb{R}^{3})$ and $L_{K_{0}}$ have a clean intersection along $\psi(K_{1})$, where $K_{1}$ is an oriented knot in $\mathbb{R}^{3}$. We may assume that for sufficiently small $\varepsilon>0$, $\psi(L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3})\subset L_{K_{0}}$. We fix an orientation of $\psi(K_{1})$ via the diffeomorphism $\left.\psi\right|_{K_{1}}\colon K_{1}\to\psi(K_{1})$ and define $a\in\mathbb{Z}$ by $\psi(K_{1})]=a[K_{0}]\in H_{1}(L_{K_{0}})$. Let $(-1)^{\sigma}$ denote the sign of orientation of the embedding $\left.\psi\right|_{L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}}\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$. Then, the claim of the proposition is equivalent to that $a\in\{1,-1\}$ and $\sigma\equiv 0\mod 2$. Since $\sigma\equiv 0\mod 2$ is already shown in [16, Lemma 4.4], we only give a proof for $a\in\{1,-1\}$ (see also Remark A.2 below).
@@
Since $\psi(\mathbb{R}^{3})$ and $L_{K_{0}}$ intersects cleanly along $\psi(K_{1})$ and $\psi(K_{1})$ is connected, we can apply [17, Theorem 3.4.11, Lemma 3.4.12] to get the following results: From the pair $(f,g)$ with $f||_{C^{1}}$ sufficiently small, we can construct a Hamiltonian $H_{f}\colon T^{*}\mathbb{R}^{3}\to\mathbb{R}$ supported in a neighborhood of $\psi(K_{1})$ and an almost complex structure $J_{g}$ on $T^{*}\mathbb{R}^{3}$ such that the pair $(H_{f},J_{g})$ is regular and the following hold:
  • $P_{H_{f}}\cap L_{K_{0}}=\{x_{0},x_{1}\}$.
  • There is a bijection
@@
since $\psi(K_{1})]\in H_{1}(L_{K_{0}})$ is represented by a loop obtained by concatenating the two paths: $\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{1}(-s)$ and $\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{2}(s)$. From this observation, the differential $d_{J_{g}}\colon\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{1}\to\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{0}$ is given by
$d_{J_{g}}(x_{1})=(\lambda^{k}+\lambda^{k+a})x_{0}.$
Therefore, $HF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};(H_{f},J_{g}))$ is computed as follow (we omit writing $(H_{f},J_{g})$):
@@
The fact $\sigma\equiv 0\mod 2$ was proved in [16, Lemma 4.4] by introducing $\mathbb{Z}$-gradings of Lagrangian submanifolds which are lifts of orientations of Lagrangian submanifolds. Analogous to [16, Remark 2.9, Lemma 4.4], by an argument on $\mathbb{Z}/2$-grading in the present case, we can deduce $\sigma\equiv 0\mod 2$ from $\sigma_{\psi}\equiv\sigma_{\operatorname{id}}\mod 2$.

### v3

@@ 3.2.3 Application of $\mathbb{Q}$-valued augmentations
Figure 1 below illustrates $u(\mathbb{R}\times[0,1])\subset T^{*}\mathbb{R}^{3}$ and the three paths in $L_{K_{0}}$. The group $\mathbb{R}$ acts on $\mathcal{M}_{(H,J),A}(x,y)$ by $(u\cdot\sigma)(s,t)=u(s+\sigma,t)$ for every $\sigma\in\mathbb{R}$.
[…]
To define the $\mathbb{Z}/2$-graded Floer cochain complex, we refer to [17, Section 2.f]. We take the pair $(H,J)$ to be regular so that the moduli space $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is cut out transversely for every $x,y\in P_{H}\cap L_{K_{0}}$ such that $x\neq y$. The dimension of $\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is equal to |x||y|1|x|-|y|-1 modulo $2$. Moreover, by the Gromov compactness, the union of 0-dimensional components of $\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is a finite set. We define $n_{(H,J),A}(x,y)\in\mathbb{F}_{2}$ to be the modulo $2$ count of the number of the 0-dimensional components of $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$.
@@
For $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ in the statement of Proposition 2.7, let us set $\psi\coloneqq\varphi^{-1}$. Then, we can change the setup as follows: $\psi(\mathbb{R}^{3})$ and $L_{K_{0}}$ have a clean intersection along $\psi(K_{1})$, where $K_{1}$ is an oriented knot in $\mathbb{R}^{3}$. We may assume that for sufficiently small $\varepsilon>0$, $\psi(L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3})\subset L_{K_{0}}$. We fix an orientation of $\psi(K_{1})$ via the diffeomorphism $\left.\psi\right|_{K_{1}}\colon K_{1}\to\psi(K_{1})$ and define $a\in\mathbb{Z}$ by $\psi(K_{1})]=a[K_{0}]\in H_{1}(L_{K_{0}})$. Let $(-1)^{\sigma}$ denote the sign of orientation of the embedding $\left.\psi\right|_{L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}}\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$. Then, the claim of the proposition is equivalent to that $a\in\{1,-1\}$ and $\sigma\equiv 0\mod 2$. Since $\sigma\equiv 0\mod 2$ is already shown in [15, Lemma 4.4], we only give a proof for $a\in\{1,-1\}$ (see also Remark A.2 below).
@@
Since $\psi(\mathbb{R}^{3})$ and $L_{K_{0}}$ intersects cleanly along $\psi(K_{1})$ and $\psi(K_{1})$ is connected, we can apply [16, Theorem 3.4.11, Lemma 3.4.12] to get the following results: From the pair $(f,g)$ with $f||_{C^{1}}$ sufficiently small, we can construct a Hamiltonian $H_{f}\colon T^{*}\mathbb{R}^{3}\to\mathbb{R}$ supported in a neighborhood of $\psi(K_{1})$ and an almost complex structure $J_{g}$ on $T^{*}\mathbb{R}^{3}$ such that the pair $(H_{f},J_{g})$ is regular and the following hold: $P_{H_{f}}\cap L_{K_{0}}=\{x_{0},x_{1}\}$ and there is a bijection
@@
since $\psi(K_{1})]\in H_{1}(L_{K_{0}})$ is represented by a loop obtained by concatenating the two paths: $\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{1}(-s)$ and $\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{2}(s)$. From this observation, the differential $d_{J_{g}}\colon\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{1}\to\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{0}$ is given by $d_{J_{g}}(x_{1})=(\lambda^{k}+\lambda^{k+a})x_{0}$. Therefore, $HF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};(H_{f},J_{g}))$ is computed as follow (we omit writing $(H_{f},J_{g})$):
@@
The fact $\sigma\equiv 0\mod 2$ was proved in [15, Lemma 4.4] by introducing $\mathbb{Z}$-gradings of Lagrangian submanifolds which are lifts of orientations of Lagrangian submanifolds. Analogous to [15, Remark 2.9, Lemma 4.4], by an argument on $\mathbb{Z}/2$-grading in the present case, we can deduce $\sigma\equiv 0\mod 2$ from $\sigma_{\psi}\equiv\sigma_{\operatorname{id}}\mod 2$.